OpenStudy (anonymous):

Quick Question on Rodrigues' formula, the formula is given with Taus sometimes and with t other times. Tau = t*|w| but i see Tau inside the sin/cos functions sometimes with |w|...what is the correct formula?

OpenStudy (anonymous):

t = tau. Just plug in tau into the rodrigues formula for t. also for Q3, assume tau is 1.

OpenStudy (anonymous):

so in the equations where tau is inside a cosine or sine with ||w||, i still have ||w|| in there right?

OpenStudy (anonymous):

\[e^{w *\tau}= 1(Identity) +\sin(||w||*\tau)\frac{ w }{ ||w|| }+(1-\cos(||w||*\tau))\frac{ w ^{2} }{ ||w||^{2} }\]

OpenStudy (anonymous):

Yes that's correct.

OpenStudy (anonymous):

ok great, thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

I calculated tau in problem 3 to be 1.1179, I should go back and change it to 1?

OpenStudy (anonymous):

Ha no don't worry about it. You're not the only person to have done that but actually that calculation you made was under the assumption that mag(omega) = 1, which you can't normally assume.

OpenStudy (anonymous):

Is there a way to do our programming to the robot arm through Virtual Lab? I wasn't sure if the com ports mapped over...

OpenStudy (anonymous):

I'm not an expert on this by I would not think the com ports would map like that. Let me know if you get it to work.